Please help me with this problem. Determine if (3/4 . -2) is a solution to
-4x - 7y =11 Please show all work.

Answer by jim_thompson5910(35256)

-4x - 7y =11


-4(3/4) - 7(-2) =11 ... plug in x=3/4 and y=-2


-3 + 14 =11


11 = 11


So the solution checks out.